The Length of a pool is 3 times its width and the pool is surrounded by a grass walk 4 feet wide. If the total area enclosed by the walk(of uniform width) is 684 square feet, find the dimensions of the pool.
Answer by checkley77(12844)

L=3W
AREA=(L+2*4)(W+2*4)
684=(L+8)(W+8)
684=(3W+8)(W+8)
684=3W^2+8W+24W+64
684=3W^2+32W+64
3W^2+32W+64-684=0
3W^2+32W-620=0
(3X+62)(W-10)=0
W-10=0
W=10 FT. ANS FOR THE WIDTH OF THE POOL.
L=3*10=30 FT. ANS. FOR THE LENGTH OF THE POOL.
PROOF:
684=(30+8)(10+8)
684=38*18
684=684